# Comparison of 4.5PN and 2SF gravitational energy fluxes from quasicircular compact binaries

### v1

@@ Abstract
The past three years have seen two significant advances in models of gravitational waveforms emitted by quasicircular compact binaries in two regimes: the weak-field, post-Newtonian regime, in which the gravitational wave energy flux has now been calculated to fourth-and-a-half post-Newtonian order (4.5PN) [Phys. Rev. Lett. 131, 121402 (2023)]; and the small-mass-ratio, gravitational self-force regime, in which the flux has now been calculated to second perturbative order in the mass ratio (2SF) [Phys. Rev. Lett. 127, 151102 (2021)]. We compare these results and find excellent agreement for the total flux, showing consistency between the two calculations at all available PN and SF orders. However, although the total fluxes agree, we find disagreements in the fluxes due to individual spherical-harmonic modes of the waveform, strongly suggesting the two waveforms might be in different asymptotic frames.
@@ I Introduction
Recently, two important milestones were achieved in this effort, specifically in perturbation methods for general relativity (GR) applied to gravitational waves from compact binary systems on quasicircular orbits: (i) the gravitational self-force (GSF) approach was pushed to second order in the small mass ratio, both for the binary’s conserved mass-energy [4] and the GW emission [5, 6]; (ii) the post-Newtonian (PN) approximation was completed at the 4PN order for the equations of motion [7, 8, 9, 10, 11, 12] and at the 4.5PN order (beyond the Einstein quadrupole formula) for the energy flux [13, 14, 15, 16, 17, 18, 19], while the dominant $(\ell,m)=(2,2)$ mode was obtained at 4PN order [18, 19]. The aim of this paper is to compare the salient results of the two approaches, confirming the preliminary agreement found in [20].
@@
We write the asymptotic gravitational wave in a transverse traceless (TT) gauge, and decompose it into a “plus” mode $h_{+}$ and a “cross” mode $h_{\times}$. We can then expand the waveform $h\equiv h_{+}-\mathrm{i}h_{\times}$ into spin-weighted spherical harmonics (with weight $-2$) as follows:
@@
where ${\cal F}_{\mathcal{H}_{i}}(x,\nu)$ is the flux of energy into the horizon of the black hole of mass $m_{i}$. Here $\hat{h}_{\ell m}$, $\Xi$, and ${\cal F}_{\mathcal{H}_{i}}$ are dimensionless and therefore can only be functions of dimensionless combinations of the binary’s frequency $\omega$ and masses $m_{i}$ — and all such combinations can be written in terms of $x$ and $\nu$. Equations (4) allow us to apply the chain rule
@@
when acting on $h_{\ell m}$ with derivatives of $u$. The asymptotic energy flux carried by GWs reads
@@
Note that $m=0$ modes do not contribute for quasicircular orbits at the orders we consider, and a factor of 2 has appeared when expressing the sum in terms of positive $m$ modes, due to the relationship $h_{\ell(-m)}=(-)^{m}{h}_{\ell m}^{*}$ (the star * indicates complex conjugation). Also note that although $\dot{h}_{\ell m}$ is a function of $(x,\nu,\psi)$, the modulus squared $|\dot{h}_{\ell m}|^{2}$ depends only on $(x,\nu)$.
These fluxes are related in a one-to-one manner to the moduli of the $h_{\ell m}$ modes, but do not carry any information about the phases. In the PN approach, which assumes small orbital frequencies, we obtain the fluxes as an expansion in the small parameter $x$ of Eq. (3), with coefficients that are functions of $\nu$:
@@
where the SF order “$(n)$” indicates the order in $\nu$ at which the field equations are solved in order to obtain the corresponding flux contribution $\mathcal{F}^{(n)}_{\ell m}$.
@@
The goal of this article is to confirm the agreement up to 4.5PN. Our central conclusion is that the 2SF and 4.5PN results agree on the total flux at each available order: i.e. $\sum_{\ell m}\mathcal{F}^{pqk}_{\ell m}$ agrees between the two methods for $p=1$ and $q\in[\![0,9]\!]$ (and $k=0$ or $1$). However, we also find a notable disagreement on the individual mode contributions $\mathcal{F}^{pqk}_{\ell m}$ beyond 3.5PN order, which precisely cancels in the sum over modes. As we discuss below, this might indicate that the PN and GSF waveforms are calculated in different asymptotic frames.
@@
where the PN remainder terms are entirely controlled with the 4PN precision [28, 19]. Here $\mathrm{I}_{L}$, $\mathrm{J}_{L}$ are functionals of the matter plus gravitation pseudo stress-energy tensor $\overline{\tau}^{\alpha\beta}$ (the overbar means the PN expansion),
@@
Once the external MPM metric is determined, we expand it to leading order in the distance between source and observer, to obtain the asymptotic waveform and the GW observables at infinity. Because harmonic coordinates asymptotically exhibit logarithms of the radial coordinate, which spoil the multipolar structure of the asymptotic waveform, we introduce a (non harmonic) “radiative” coordinate system adapted to the fall-off of the metric at infinity. By radiative coordinate system we mean one such that the retarded time $u=t-r$ is a null (or asymptotically null) coordinate, thus avoiding the logarithmic deviation of the retarded time in harmonic coordinates with respect to the true light cone. Alternatively, a different MPM construction defined in [24], corrects the coordinate light-cones at every PM order so as to iteratively build the null coordinate $u$, with the far zone logarithms therefore automatically cancelled. This construction has played a crucial role in our calculation of the 4PN flux.
@@
×$\qquad\qquad\qquad\times\left[\ln\left(\frac{u-\tau}{b_{0}}\right)+\frac{11}{12}\right]+\mathcal{O}(5)\,,$ (14)
where $\mathrm{M}$ is the monopole of the multipole expansion (i.e. the Arnowitt-Deser-Misner mass of the spacetime) and $b_{0}$ is an arbitrary constant length scale linked to the definition of radiative coordinates. Beyond the tail term in (II), one finds the nonlinear memory effect at orders 2.5PN and 3.5PN; the tail-of-tail at 3PN order (a cubic interaction between two masses $\mathrm{M}$ and the time-varying quadrupole); the tails-of-memory at 4PN order, which is a cubic interaction between $\mathrm{M}$ and two varying quadrupoles [17]; the spin quadrupole tail at 4PN order (interaction $\mathrm{M}\times\mathrm{S}_{i}\times\mathrm{M}_{L}$); and at 4.5PN order the quartic tail-of-tail-of-tail interation, composed of three masses $\mathrm{M}$ and the quadruole [14].
The PN-MPM formalism is applied to compact binary sources. The compact objects are modeled by point particles (we neglect spins and other finite size effects). Furthermore, in a first stage of the calculation the two masses $m_{i}$ are considered to be constant, i.e. we neglect the black hole absorption, which should be added separately at the end of the PN calculation.
@@
where ${\cal F}$ denotes the energy flux at infinity. Therefore, when solving Eq. (17) to obtain the expression of the frequency evolution (or chirp), i.e. $\dot{x}$ expressed in terms of $x$, we shall obtain extra 4PN corrections due to the horizon fluxes. Those will affect the equations of motion at 4PN order beyond the dominant 2.5PN radiation reaction, i.e. corresponding to 6.5PN radiation reaction terms. When obtaining the 4.5PN flux at infinity, Eq. (II) below, we only required the expression of the frequency chirp at 1.5PN order beyond the dominant order (see (5.3b) of [19]), hence the above 6.5PN radiation reaction terms can be safely discarded in our calculation. Our conclusion is that the BH absorption plays no role in the derivation of the energy flux at infinity (II) at this order, so we shall directly compare it with the GSF result in Sec. IV.
The equations of motion of the compact binary system have already been obtained at 4PN order [7, 8, 9, 10, 11, 12] and we extensively use the result. The most important task is the computation of the source multipole moments $\mathrm{I}_{L}$, $\mathrm{J}_{L}$ for the binary source, see Eq. (13). The source moments represent the seed of the full construction through the steps (12)–(II). The 4PN mass quadrupole moment $\mathrm{I}_{ij}$ has been obtained in [13, 15, 16]. Previous works had determined that UV divergences, due to the modelling of compact objects by point masses, appear at the 3PN order and require the use of dimensional regularization. After regularization a shift of the particle’s world lines permits to absorb the divergences. The works [13, 15, 16] showed that at the 4PN order IR divergences appear as well. For reasons explained in [9, 10], we have used a variant of dimensional regularization called the “$B\epsilon$” regularization, where the finite part at $B=0$ in Eq. (13) is applied on the top of the calculation performed in $d=3+\epsilon$ dimensions. An important point is that the regularization constant $r_{0}$ in (13) is finally cancelled in the radiative moment (II), due notably to the $r_{0}$ dependence of the tails-of-memory. Besides the 4PN mass quadrupole moment, we also need the 3PN mass octupole $\mathrm{I}_{ijk}$ and 3PN current quadrupole $\mathrm{J}_{ij}$, which have been obtained in [31] and [32], respectively.
It is known that the GW half-phase $\psi$, defined by the decomposition of Eq. (2), and the orbital phase $\phi$, which is used in the computation to parameterize the circular motion of the binary, differ by a logarithmic phase modulation at order 1.5PN, which is due to the propagation of tails in the wave zone [33, 34, 35],
@@
where $y_{0}=(m_{\text{tot}}\,\omega_{0})^{2/3}$ and $\mathcal{O}(y^{5})=\mathcal{O}(10)$. When expressing physical results in terms of the GW observables $\psi$ and $\omega$, the arbitrary scale $b_{0}$ is canceled out (see [19] for details).
@@ III Review of 2SF fluxes
where $h^{(n)}_{\ell m}$ is obtained from the coefficient of $1/r$ in the large-$r$ limit of the metric perturbation $h^{(n)}_{\alpha\beta}$ in a Bondi-Sachs (radiative) gauge [42]. In practice, we solve the Einstein equations for $h^{(n)}_{\alpha\beta}$ in the Lorenz gauge, [^2] Here the Lorenz gauge condition is $\mathring{g}^{\beta\gamma}\mathring{\nabla}_{\beta}\bar{h}_{\alpha\gamma}=0$, where $\mathring{g}^{\beta\gamma}$ is the inverse of the background metric $\mathring{g}_{\beta\gamma}$, ̊$\mathring{\nabla}_{\beta}$ is the covariant derivative compatible with $\mathring{g}_{\alpha\beta}$, and $\bar{h}_{\alpha\beta}\equiv(\mathring{g}_{\alpha}{}^{\mu}\mathring{g}_{\beta}{}^{\nu}-\frac{1}{2}\mathring{g}_{\alpha\beta}\mathring{g}^{\mu\nu})(g_{\mu\nu}-\mathring{g}_{\mu\nu})$. This gauge shares many essential features with the harmonic coordinate condition $g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}=0$ used in PN theory. which is not well behaved at large $r$ [43, 37, 38] (as in the PN case in Sec. II), and then transform to a Bondi-Sachs gauge to eliminate the ill-behaved pieces of the Lorenz-gauge solution.
@@
Note that the dependence on $\delta m_{1}$ cancels on the right-hand side of Eq. (31b), leaving only a dependence on the frequency variable $y$ [^3] We emphasize that this only represents the elimination of the split of the physical, evolving mass $m_{1}$ into a constant background mass and an evolving correction. The mass’s evolution enters directly in the waveform’s time dependence via Eq. (25c). The rate of change of the mass also enters the $\ell=0$ field equations for $h^{(2)}_{\alpha\beta}$ when time derivatives act on the metric perturbation (24[37], but the resulting contribution to the $\ell=0$ mode only enters the $\mathcal{O}(\nu^{4})$ flux. Finally, the horizon absorption enters into the field equations and the $\mathcal{O}(\nu^{3})$ flux to infinity via $F^{\Omega}_{0}$ in Eq. (26), but $F^{\Omega}_{0}$ can equivalently be calculated from the local first-order dissipative self-force rather than from fluxes.
@@ IV Comparison
Figure 1: Comparison between GSF and PN for the 2SF $\mathcal{O}(\nu)$ contribution to the (Newtonian-normalized) total flux (43). The 2SF results are shown with (blue) dots. The PN results are shown with coloured curves with integer PN orders shown with solid curves and corresponding half-integer orders shown with dashed curves.
Figure 2: Detailed comparison between the $\mathcal{O}(\nu)$ PN and GSF results for the (Newtonian-normalized) total flux. The 2SF data are shown by (blue) dots and the PN series are plotted with solid curves. As ever higher-order PN series are subtracted from the GSF results the residual at small $x$ behaves as expected. After subtracting the 3.5PN series from the GSF data, one gets the (yellow) squares which follow the $\mathcal{O}(x^{4})$ behaviour of the 4PN series (yellow curve). Subtracting the 4PN series from the SF data, one gets the (green) diamonds. This agrees with the $\mathcal{O}(x^{9/2})$ term in the 4.5PN series (green curve) though for $x\lesssim 0.06$ the numerical error in the GSF calculation dominates the comparison. After subtracting all known terms in the PN series there is evidence that the residual has the expected $x^{5}$ behaviour (dashed, red line).
Figure 3: (Left Panel) Comparison between the (Newtonian-normalized) GSF and PN total flux at different orders in the mass ratio expansion for $q=1$ ($\nu=0.25$). It is interesting to note that the 4.5PN expansion, including all orders in $\nu$ (black, dashed curve), is closely approximated by the expansion truncated at $\mathcal{O}(\nu^{2})$ (yellow, solid curve). In gray we also show the flux computed from an NR simulation by the SXS collaboration [45], specifically SXS:BBH:1132 [46]. The shaded width of the gray curve is an estimate on the numerical error in the simulation. (Right Panel) The same as the left panel but for $q=10$ ($\nu=0.0826$). The NR simulation in this plot is SXS:BBH:1107 [47].
@@
We present the results of our extensive comparison between the numerical 2SF data and the analytic 4.5PN flux given by Eq. (II). In computing the total 2SF flux, we sum the $(\ell,m)$ modes up to $\ell=7$. We find that the contribution from modes with higher $\ell$ does not affect the comparison. For the purpose of the GSF and PN comparison, we define the Newtonian-normalized flux as
^$\hat{\mathcal{F}}\equiv\frac{\mathcal{F}}{\mathcal{F}_{\rm N}}\,,\qquad\hat{\mathcal{F}}_{\ell m}\equiv\frac{\mathcal{F}_{\ell m}}{\mathcal{F}_{\rm N}}\,,$ (43)

### v2

@@ Abstract
Recent years have seen significant advances in models of gravitational waveforms emitted by quasicircular compact binaries in two regimes: the weak-field, post-Newtonian regime, in which the gravitational wave energy flux has now been calculated to fourth-and-a-half post-Newtonian order (4.5PN) [Phys. Rev. Lett. 131, 121402 (2023)]; and the small-mass-ratio, gravitational self-force regime, in which the flux has now been calculated to second perturbative order in the mass ratio (2SF) [Phys. Rev. Lett. 127, 151102 (2021)]. We compare these results and find agreement, showing consistency between the two (very distinct though both first-principle) perturbative calculations.
@@ I Introduction
Recently, two important milestones were achieved in this effort, specifically in perturbation methods for general relativity (GR) applied to gravitational waves from compact binary systems on quasicircular orbits: (i) the gravitational self-force (GSF) approach was pushed to second order in the small mass ratio, both for the binary’s conserved mass-energy [4] and the GW emission [5, 6]; (ii) the post-Newtonian (PN) approximation was completed at 4PN order for the equations of motion [7, 8, 9, 10, 11, 12] and at 4.5PN order (beyond the Einstein quadrupole formula) for the energy flux [13, 14, 15, 16, 17, 18, 19], while the dominant $(\ell,m)=(2,2)$ mode of the waveform was obtained at 4PN order [18, 19]. The aim of this paper is to compare the salient results of the two approaches, confirming the preliminary agreement found in [20].
@@
We write the asymptotic gravitational waveform in a transverse traceless (TT) gauge, and decompose it into a “plus” mode $h_{+}$ and a “cross” mode $h_{\times}$. We can then expand the waveform $h\equiv h_{+}-\mathrm{i}h_{\times}$ into spin-weighted spherical harmonics (with weight 2-2) as follows:
@@
where ${\cal F}_{\mathcal{H}_{i}}(x,\nu)$ is the flux of energy into the horizon of the black hole of mass $m_{i}$. Here $\hat{h}_{\ell m}$, ${\cal F}_{\mathcal{H}_{i}}$, and the forcing function $\Xi$ are dimensionless and therefore can only be functions of dimensionless combinations of the binary’s frequency $\omega$ and masses $m_{i}$ — and all such combinations can be written in terms of $x$ and $\nu$. Equations (4) allow us to apply the chain rule
@@
when acting on $h_{\ell m}$ with derivatives of $u$.
The asymptotic energy flux carried by GWs is
@@
Note that $m=0$ modes do not contribute for quasicircular orbits at the orders we consider, and a factor of 2 has appeared when expressing the sum in terms of positive $m$ modes, due to the relationship $h_{\ell(-m)}=(-)^{m}{h}_{\ell m}^{*}$ (the star * indicates complex conjugation). Also note that although $\dot{h}_{\ell m}$ is a function of $(x,\nu,\psi)$, the modulus squared $|\dot{h}_{\ell m}|^{2}$ depends only on $(x,\nu)$; these fluxes are related in a one-to-one manner to the moduli of the $h_{\ell m}$ modes, but do not carry any information about the phases.
In the PN approach, which assumes small orbital frequencies, we obtain the fluxes as an expansion in the small parameter $x$ of Eq. (3), with coefficients that are functions of $\nu$:
@@
where the SF order “(n)(n)” indicates the order in $\nu$ at which the field equations are solved in order to obtain the corresponding flux contribution $\mathcal{F}^{(n)}_{\ell m}$.
@@
The goal of this article is to confirm agreement up to 4.5PN. Our central conclusion is that the 2SF and 4.5PN results are consistent with each other at each available order: i.e. $\sum_{\ell m}\mathcal{F}^{pqk}_{\ell m}$ agrees between the two methods for $p=1$ and $q\in[\![0,9]\!]$ (and $k=0$ or $1$). However, at 4.5PN in particular, the comparison is hampered by poor overlap of the regions of validity of the two approximations: the 4.5PN flux is only in its asymptotic regime of validity for small values of $x$ in which numerical error in the 2SF data becomes significant enough that it interferes with the comparison.
@@
where the PN remainder terms are entirely controlled with the 4PN precision [28, 19]. Here $\mathrm{I}_{L}$, $\mathrm{J}_{L}$ are functionals of the matter plus the gravitation stress-energy pseudo-tensor $\overline{\tau}^{\alpha\beta}$ (the overbar means the PN expansion),
@@
Once the external MPM metric is determined, we expand it to leading order in the distance between source and observer, to obtain the asymptotic waveform and the GW observables at infinity. Because harmonic coordinates asymptotically exhibit logarithms of the radial coordinate, which spoil the multipolar structure of the asymptotic waveform, we introduce a (non harmonic) “radiative” coordinate system adapted to the fall-off of the metric at infinity. By radiative coordinate system we mean one such that the retarded time u=tru=t-r is a null (or asymptotically null) coordinate, thus avoiding the logarithmic deviation of the retarded time in harmonic coordinates with respect to the true light cone. Alternatively, a different MPM construction defined in [24], corrects the coordinate light-cones at every PM order so as to iteratively build the null coordinate $u$, with the far zone logarithms therefore automatically cancelled. This construction has played a crucial role in our calculation of the 4PN flux.
@@
$\qquad\qquad\qquad\times\left[\ln\left(\frac{u-\tau}{2b_{0}}\right)+\frac{11}{12}\right]+\mathcal{O}(5)\,$, (14)
where $\mathrm{M}$ is the monopole of the multipole expansion (i.e. the Arnowitt-Deser-Misner mass of the spacetime) and $b_{0}$ is an arbitrary constant length scale linked to the definition of radiative coordinates. Beyond the tail term in (II), one finds the nonlinear memory effect at orders 2.5PN and 3.5PN; the tail-of-tail at 3PN order (a cubic interaction between two masses $\mathrm{M}$ and the time-varying quadrupole); the tails-of-memory at 4PN order, which is a cubic interaction between $\mathrm{M}$ and two varying quadrupoles [17]; the spin quadrupole tail at 4PN order (interaction $\mathrm{M}\times\mathrm{S}_{i}\times\mathrm{M}_{L}$); and at 4.5PN order the quartic tail-of-tail-of-tail interaction, composed of three masses $\mathrm{M}$ and the quadrupole [14].
The PN-MPM formalism is applied to compact binary sources. The compact objects are modelled by point particles (we neglect spins and other finite size effects). Furthermore, in a first stage of the calculation the two masses $m_{i}$ are considered to be constant, i.e. we neglect the black hole absorption, which should be added separately at the end of the PN calculation.
@@
[…]
The equations of motion of the compact binary system have already been obtained at 4PN order [7, 8, 9, 10, 11, 12] and we extensively use the result. The most important task is the computation of the source multipole moments $\mathrm{I}_{L}$, $\mathrm{J}_{L}$ for the binary source, see Eq. (13). The source moments represent the seed of the full construction through the steps (12)–(II). The 4PN mass quadrupole moment $\mathrm{I}_{ij}$ has been obtained in [13, 15, 16]. Previous works had determined that UV divergences, due to the modelling of compact objects by point masses, appear at the 3PN order and require the use of dimensional regularization. After regularization a shift of the particle’s world lines permits to absorb the divergences. The works [13, 15, 16] showed that at the 4PN order IR divergences appear as well. For reasons explained in [9, 10], we have used a variant of dimensional regularization called the $B\epsilon$” regularization, where the finite part at $B=0$ in Eq. (13) is applied on the top of the calculation performed in $d=3+\epsilon$ dimensions. An important point is that the regularization constant $r_{0}$ in (13) is finally cancelled in the radiative moment (II), due notably to the $r_{0}$ dependence of the tails-of-memory. Besides the 4PN mass quadrupole moment, we also need the 3PN mass octupole $\mathrm{I}_{ijk}$ and 3PN current quadrupole $\mathrm{J}_{ij}$, which have been obtained in [31] and [32], respectively.
It is known that the GW half-phase $\psi$, defined by the decomposition of Eq. (2), and the orbital phase $\phi$, which is used in the computation to parametrize the circular motion of the binary, differ by a logarithmic phase modulation at order 1.5PN, which is due to the propagation of tails in the wave zone [33, 34, 35],
@@
where $y_{0}=(m_{\text{tot}}\,\omega_{0})^{2/3}$ and $\mathcal{O}(y^{5})=\mathcal{O}(10)$. When expressing physical results in terms of the GW observables $\psi$ and $\omega$, the arbitrary scale $b_{0}$ is cancelled out (see [19, 36] for details).
@@ III Review of 2SF fluxes
where $h^{(n)}_{\ell m}$ is obtained from the coefficient of $1/r$ in the large-$r$ limit of the metric perturbation $h^{(n)}_{\alpha\beta}$ in a Bondi-Sachs (radiative) gauge [43]. In practice, we solve the Einstein equations for $h^{(n)}_{\alpha\beta}$ in the Lorenz gauge, [^2] Here the Lorenz gauge condition is $\mathring{g}^{\beta\gamma}\mathring{\nabla}_{\beta}\bar{h}_{\alpha\gamma}=0$, where $\mathring{g}^{\beta\gamma}$ is the inverse of the background metric $\mathring{g}_{\beta\gamma}$, $\mathring{\nabla}_{\beta}$ is the covariant derivative compatible with $\mathring{g}_{\alpha\beta}$, and $\bar{h}_{\alpha\beta}\equiv(\mathring{g}_{\alpha}{}^{\mu}\mathring{g}_{\beta}{}^{\nu}-\frac{1}{2}\mathring{g}_{\alpha\beta}\mathring{g}^{\mu\nu})(g_{\mu\nu}-\mathring{g}_{\mu\nu})$. This gauge shares many essential features with the harmonic coordinate condition $g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}=0$ used in PN theory. which is not well behaved at large $r$ [44, 38, 39] (as in the PN case in Sec. II), and then transform to a Bondi-Sachs gauge to eliminate the ill-behaved pieces of the Lorenz-gauge solution.
The transformation to the Bondi-Sachs gauge is explained in Ref. [45]. In that reference, we also discovered an unforeseen contribution to $h^{(2)}_{\ell m}$, due to the interaction between oscillatory modes and gravitational-memory modes in $h^{(1)}_{\alpha\beta}$. The new “memory distortion” term contributes to the flux ${\cal F}^{(2)}$, but we do not include it in our comparisons in this paper because (i) numerical results for it have only recently been obtained and are not yet published, and (ii) the contribution is numerically very small and appears to enter only at 5PN [46].
@@
Note that the dependence on $\delta m_{1}$ cancels on the right-hand side of Eq. (31b), leaving only a dependence on the frequency variable yy. [^3] We emphasize that this only represents the elimination of the split of the physical, evolving mass $m_{1}$ into a constant background mass and an evolving correction. The mass’s evolution enters directly in the waveform’s time dependence via Eq. (25c). The rate of change of the mass also enters the $\ell=0$ field equations for $h^{(2)}_{\alpha\beta}$ when time derivatives act on the metric perturbation (24[38], but the resulting contribution to the $\ell=0$ mode only enters the $\mathcal{O}(\nu^{4})$ flux. Finally, the horizon absorption enters into the field equations and the $\mathcal{O}(\nu^{3})$ flux to infinity via $F^{\Omega}_{0}$ in Eq. (26), but $F^{\Omega}_{0}$ can equivalently be calculated from the local first-order dissipative self-force rather than from fluxes.
@@
$A^{(2)}_{\ell m}$ $=\frac{1}{2\bigl|\tilde{h}^{(1)}_{\ell m}\bigr|}\biggl[\tilde{h}^{(1)*}_{\ell m}\biggl(\tilde{h}^{(2)}_{\ell m}-\frac{2m_{\mathrm{tot}}\omega^{(1)}}{3x^{1/2}}\partial_{x}\tilde{h}^{(1)}_{\ell m}\biggr)+\text{c.c.}\biggr]\,.$ (40b)
@@ IV Comparison
Figure 1: Comparison between GSF and PN for the 2SF $\mathcal{O}(\nu)$ contribution to the (Newtonian-normalized) total flux (43). The 2SF results are shown with (blue) dots. The PN results are shown with coloured curves; integer PN orders are shown with solid curves and corresponding half-integer orders are shown with dashed curves.
Figure 2: Detailed comparison of $\hat{\mathcal{F}}^{(2)}$, the $\mathcal{O}(\nu)$ coefficient of the (Newtonian-normalized) total flux, computed from (i) numerical GSF and (ii) PN theory. The 2SF data are shown by (blue) dots and the corresponding 3.5PN series is plotted as a solid (blue) curve. As ever higher-order PN series are subtracted from the 2SF results, the residuals are compared with the next term in the PN series at small $x$. After subtracting the 3.5PN series from the 2SF data, one gets the (yellow) squares which follow the 4PN term (yellow curve). Subtracting the 4PN series from the SF data, one gets the (green) diamonds, which are compared against the 4.5PN term (green curve). The amplitude of the residual with the 4PN series is consistent with the amplitude of the 4.5PN term, although the slope of the residual is unclear, partly due to numerical noise. Further subtracting the 4.5PN series gives the (red) triangles. For sufficiently small values of $x$, we would expect the residual to scale as $x^{5}$. We plot a (red) dashed reference $x^{5}$ curve here but the accuracy of our numerical data is insufficient to claim agreement. One reason that the agreement with 4.5PN is not clear can be seen by considering the dominant $(2,2)$ mode: over the frequency range of our numerical GSF data there is evidence that the residual with 4PN has not reached the asymptotic regime – see Fig. 3. The agreement between PN and GSF results is clearer for some individual $(\ell,m)$ modes – see Figs. 3, 5 and the figures in Appendix B.
Figure 3: (Top) Detailed comparison of $\hat{\mathcal{F}}^{(2)}_{22}$, the $\mathcal{O}(\nu)$ coefficient of the $(\ell,m)=(2,2)$ mode of the (Newtonian-normalized) flux, computed from (i) numerical GSF and (ii) PN theory. The conventions and colorings are the same as in Fig. 2, but the blue curve now shows the 3PN series and the additional upside-down (purple) triangles show the residual after subtracting the 3PN series from the 2SF data. The latter residual closely follows the 3.5PN term (purple curve). The results of comparisons with higher PN orders are more ambiguous due to numerical noise in the GSF data. For example, the residual after subtracting the 4PN series from the 2SF result (green diamonds) appears at first sight to approach the 4.5PN term [before degrading again for even smaller values of $x$, due to numerical noise], but this is in fact an illusion caused by plotting the absolute magnitude of the residual. (Bottom) A subset of the above data (without taking the absolute magnitude) on a linear-log scale. Over the range of frequency values where we have numerical data, the residual between the GSF and 4PN is negative, but the 4.5PN term is positive. Thus, in order for these to agree asymptotically, there must be a zero crossing in the residual data at smaller values of $x$ than we have access to. Only after this zero crossing would we expect the residual and the 4.5PN curve to agree. The agreement between GSF and PN for other $(\ell,m)$ modes is clearer – see Appendix B.
Figure 4: (Left panel) Detailed comparison of $\hat{\mathcal{F}}^{(2)}_{33}$, the $\mathcal{O}(\nu)$ coefficient of the $(3,3)$ mode of the (Newtonian-normalized) flux, computed from (i) numerical GSF and (ii) PN theory. The residuals between the GSF data and the various PN series (dotted lines) do not clearly follow the next term in the PN series (full lines) in this frequency range. However, the relative difference between the dotted lines and the corresponding full lines become smaller as one goes to smaller frequencies. This suggests that at smaller frequencies, one could credibly reach the asymptotic regime where the two curves overlap. Finally, the residual with the 4.5PN term approaches a 5PN slope, as shown by the $\mathcal{O}(x^{5})$ reference line. (Right panel) Same as the left panel but for the (3,2) mode. The residual with the 3PN series (purple triangles) agrees perfectly with the 3.5PN term (purple line). The residual with the 3.5PN terms (orange squares) undergoes a passage through zero, then becomes noisy for $x<0.05$, so it is not possible to claim agreement with the 4PN term (yellow line). Nonetheless, when going to the next order, we find that the residual with the 4PN series (green diamonds) agrees fairly well with the 4.5PN term (green line). Although the asymptotic regime is not yet reached, the relative error between the two curves decreases for smaller frequencies, and suggests better agreement deeper in the weak field regime. Finally, the residual with the 4.5PN term exhibits a very clear 5PN slope, as shown by the perfect fit with a $\mathcal{O}(x^{5})$ reference line. [^5] There should be also some logarithmic terms $\mathcal{O}(x^{5}\ln x)$, but one cannot distinguish the slopes of $x^{5}$ and $x^{5}\ln x$ on the scale of the plot.
@@
We present the results of our extensive comparison between the numerical 2SF data and the analytic 4.5PN flux given by Eq. (II). In computing the total 2SF flux, we sum the $(\ell,m)$ modes up to $\ell=6$ (the contribution from modes with higher $\ell$ is numerically small and does not substantially affect the comparison). For the purpose of the GSF and PN comparison, we define the Newtonian-normalized flux as
$\hat{\mathcal{F}}\equiv\frac{\mathcal{F}}{\mathcal{F}_{\rm N}}\,,\qquad\hat{\mathcal{F}}_{\ell m}\equiv\frac{\mathcal{F}_{\ell m}}{\mathcal{F}_{\rm N}}\,$, (43)